# LLM-based feature generation from text for interpretable machine learning

### v1

@@ 3.1 Computational resources and complexity
The non-finetuned LLama2 13B GPTQ Chat model was run locally with the use of GPU computation as the crucial part of the experiment. The GPU used was NVIDIA RTX A4000 with 16GB of video memory, all of which were utilized for one instance of the LLama2 model.
The time complexity was around 40-60 minutes per feature for 3000 abstracts, making the total runtime of the experiment around 60 hours for CORD-19 and 40 hours for M17+.
@@ 3.2 LLM-augmented datasets
The features generated by the language model were added to both CORD-19 and M17+ datasets, resulting in the addition of 62 columns. Features rigor, novelty, grammar, accessibility, and replicability were ordinally encoded while disciplines, research were dummy-encoded, increasing the number of columns.
After LLM inference of features, only 2 articles in the CORD-19 dataset were assigned feature values outside of the allowed feature space. These instances were removed, making the final length of CORD-19 2,998 articles (instances). All LLM-generated feature values for the M17+ articles were valid, resulting in a dataset of 2,000 instances. The first few rows from both generated datasets are included in Table 4 and Table 5.
@@
Table 4: Sample of enriched CORD-19 dataset with added LLM-generated features. Target is a binary variable - with 1 representing works cited more than the median value adjusted in time, 0 represents articles with lower or equal to the median citation count (details in [4]).
@@
Table 5: Sample of enriched M17+ dataset with added LLM-generated features. Target is an ordinal variable of human evaluation of the quality of articles, 1 indicates works of world-class quality, 5 means mediocre articles of low importance and impact.
3.3 Analysis of relationship between target and LLM-features
@@
To formally test whether generated features have a relationship with the respective target variables, we employed the chi-squared test of independence at significance level $\alpha=0.05$.
The results are shown in Table 6 for CORD-19 and in Table 7 for M17+ datasets.
Table 6: Statistically significant features according to the p-value of the relation between target and LLM-generated features, CORD-19. The star notation indicates the level of statistical significance based on p-values: ${}^{***}p<0.001$, ${}^{**}p<0.01$, and ${}^{*}p<0.05$. 31 out of 62 generated features are statistically significant for the prediction of citation rate.
Feature Importance p-value Feature Importance p-value
Rigor [^1] Categorical features based on abstract quality and contents. *** $1.62\times 10^{-7}$ Nano-tech [^2] Binary features belonging to disciplines values. * $1.08\times 10^{-2}$
Novelty [^1] Categorical features based on abstract quality and contents. ** $8.59\times 10^{-4}$ Basic Medicine [^2] Binary features belonging to disciplines values. *** $3.26\times 10^{-17}$
Math [^2] Binary features belonging to disciplines values. ** $5.44\times 10^{-3}$ Clinical Medicine [^2] Binary features belonging to disciplines values. *** $5.16\times 10^{-5}$
CompSci [^2] Binary features belonging to disciplines values. *** $4.87\times 10^{-6}$ HealthSci [^2] Binary features belonging to disciplines values. ** $3.65\times 10^{-3}$
PhysSci [^2] Binary features belonging to disciplines values. *** $7.98\times 10^{-12}$ MedBiotech [^2] Binary features belonging to disciplines values. *** $1.21\times 10^{-6}$
ChemSci [^2] Binary features belonging to disciplines values. *** $1.03\times 10^{-14}$ OtherMedSci [^2] Binary features belonging to disciplines values. *** $9.75\times 10^{-7}$
EarthSci [^2] Binary features belonging to disciplines values. *** $4.85\times 10^{-8}$ Agriculture [^2] Binary features belonging to disciplines values. *** $5.78\times 10^{-9}$
BioSci [^2] Binary features belonging to disciplines values. *** $2.31\times 10^{-15}$ AnimalSci [^2] Binary features belonging to disciplines values. *** $4.49\times 10^{-11}$
OtherNatSci [^2] Binary features belonging to disciplines values. *** $1.99\times 10^{-8}$ VeterinarySci [^2] Binary features belonging to disciplines values. *** $8.40\times 10^{-13}$
CivEng [^2] Binary features belonging to disciplines values. * $2.03\times 10^{-2}$ AgriculBiotech [^2] Binary features belonging to disciplines values. ** $3.37\times 10^{-3}$
MechEng [^2] Binary features belonging to disciplines values. *** $6.33\times 10^{-6}$ OtherAgriculSci [^2] Binary features belonging to disciplines values. ** $4.30\times 10^{-4}$
ChemEng [^2] Binary features belonging to disciplines values. *** $9.47\times 10^{-9}$ Psychology [^2] Binary features belonging to disciplines values. ** $1.87\times 10^{-3}$
MatEng [^2] Binary features belonging to disciplines values. * $1.05\times 10^{-2}$ Law [^2] Binary features belonging to disciplines values. * $1.42\times 10^{-2}$
EnviroEng [^2] Binary features belonging to disciplines values. ** $4.42\times 10^{-3}$ Descriptive [^3] Binary features belonging to research values. * $1.79\times 10^{-2}$
EnviroBiotech [^2] Binary features belonging to disciplines values. *** $5.41\times 10^{-7}$ Questionnaires [^3] Binary features belonging to research values. * $1.59\times 10^{-2}$
IndustBiotech [^2] Binary features belonging to disciplines values. *** $8.33\times 10^{-9}$
@@
Table 7: Statistically significant features according to the p-value of the relation between target and LLM-generated features, M17+. The star notation indicates the level of statistical significance based on p-values: ${}^{***}p<0.001$, ${}^{**}p<0.01$, and ${}^{*}p<0.05$. 58 out of 62 generated features are statistically significant for the prediction of evaluation grade.
Feature Importance p-value Feature Importance p-value
Rigor [^1] Categorical features based on abstract quality and contents. *** $7.08\times 10^{-71}$ VeterinarySci [^2] Binary features belonging to disciplines values. *** $1.21\times 10^{-25}$
Novelty [^1] Categorical features based on abstract quality and contents. *** $1.99\times 10^{-47}$ AgriculBiotech [^2] Binary features belonging to disciplines values. *** $1.76\times 10^{-8}$
Grammar [^1] Categorical features based on abstract quality and contents. *** $4.69\times 10^{-20}$ OtherAgriculBiotech [^2] Binary features belonging to disciplines values. *** $1.64\times 10^{-23}$
@@
Replicability [^1] Categorical features based on abstract quality and contents. *** $7.97\times 10^{-44}$ Psychology [^2] Binary features belonging to disciplines values. *** $3.73\times 10^{-25}$
Math [^2] Binary features belonging to disciplines values. *** $7.00\times 10^{-16}$ EconomicsBusiness [^2] Binary features belonging to disciplines values. *** $1.81\times 10^{-5}$
CompSci [^2] Binary features belonging to disciplines values. *** $8.25\times 10^{-33}$ Education [^2] Binary features belonging to disciplines values. *** $1.05\times 10^{-6}$
PhysSci [^2] Binary features belonging to disciplines values. *** $3.17\times 10^{-39}$ Sociology [^2] Binary features belonging to disciplines values. *** $2.23\times 10^{-11}$
ChemSci [^2] Binary features belonging to disciplines values. *** $2.04\times 10^{-28}$ Law [^2] Binary features belonging to disciplines values. *** $8.33\times 10^{-12}$
EarthSci [^2] Binary features belonging to disciplines values. *** $1.36\times 10^{-24}$ PoliticalSci [^2] Binary features belonging to disciplines values. ** $4.53\times 10^{-3}$
BioSci [^2] Binary features belonging to disciplines values. *** $1.68\times 10^{-51}$ Social Econ Geography [^2] Binary features belonging to disciplines values. *** $6.61\times 10^{-7}$
OtherNatSci [^2] Binary features belonging to disciplines values. *** $4.43\times 10^{-36}$ MediaCommunications [^2] Binary features belonging to disciplines values. *** $2.07\times 10^{-11}$
CivEng [^2] Binary features belonging to disciplines values. *** $2.08\times 10^{-6}$ OtherSocSci [^2] Binary features belonging to disciplines values. *** $8.70\times 10^{-19}$
ElectrEng [^2] Binary features belonging to disciplines values. *** $1.17\times 10^{-11}$ HistoryArchaelogy [^2] Binary features belonging to disciplines values. *** $6.29\times 10^{-13}$
MechEng [^2] Binary features belonging to disciplines values. *** $6.88\times 10^{-6}$ LanguagesLiterature [^2] Binary features belonging to disciplines values. *** $3.62\times 10^{-13}$
ChemEng [^2] Binary features belonging to disciplines values. *** $1.50\times 10^{-9}$ PhilosophyEthicsReligion [^2] Binary features belonging to disciplines values. *** $2.32\times 10^{-10}$
MatEng [^2] Binary features belonging to disciplines values. *** $3.84\times 10^{-5}$ Arts [^2] Binary features belonging to disciplines values. ** $1.09\times 10^{-3}$
EnviroEng [^2] Binary features belonging to disciplines values. *** $4.70\times 10^{-11}$ OtherArts [^2] Binary features belonging to disciplines values. *** $7.49\times 10^{-13}$
EnviroBiotech [^2] Binary features belonging to disciplines values. *** $1.53\times 10^{-18}$ Applied [^3] Binary features belonging to research values. *** $7.52\times 10^{-10}$
IndustBiotech [^2] Binary features belonging to disciplines values. *** $5.26\times 10^{-15}$ Conclusive [^3] Binary features belonging to research values. *** $5.23\times 10^{-7}$
Nano-tech [^2] Binary features belonging to disciplines values. *** $8.65\times 10^{-8}$ Correlational [^3] Binary features belonging to research values. ** $2.80\times 10^{-3}$
OtherEngTech [^2] Binary features belonging to disciplines values. *** $1.15\times 10^{-11}$ Descriptive [^3] Binary features belonging to research values. ** $1.02\times 10^{-2}$
@@
Basic Medicine [^2] Binary features belonging to disciplines values. *** $1.33\times 10^{-41}$ Ethnographic [^3] Binary features belonging to research values. * $4.52\times 10^{-2}$
@@
Clinical Medicine [^2] Binary features belonging to disciplines values. *** $1.60\times 10^{-21}$ Experiment [^3] Binary features belonging to research values. ** $1.46\times 10^{-3}$
HealthSci [^2] Binary features belonging to disciplines values. *** $7.52\times 10^{-18}$ Exploratory [^3] Binary features belonging to research values. *** $1.02\times 10^{-6}$
MedBiotech [^2] Binary features belonging to disciplines values. *** $1.31\times 10^{-27}$ Focus groups [^3] Binary features belonging to research values. *** $1.18\times 10^{-7}$
OtherMedSci [^2] Binary features belonging to disciplines values. *** $2.52\times 10^{-29}$ Interviews [^3] Binary features belonging to research values. *** $4.81\times 10^{-5}$
@@
Agriculture [^2] Binary features belonging to disciplines values. *** $3.74\times 10^{-17}$ Mixed methods [^3] Binary features belonging to research values. *** $3.51\times 10^{-14}$
AnimalSci [^2] Binary features belonging to disciplines values. *** $5.21\times 10^{-26}$ Qualitative [^3] Binary features belonging to research values. * $1.62\times 10^{-2}$
Quantitative [^3] Binary features belonging to research values. *** $2.43\times 10^{-4}$ Questionnaires [^3] Binary features belonging to research values. *** $1.52\times 10^{-8}$
The M17+ dataset has a larger number of features that indicate a statistically significant relationship between the LLM-generated features and the awarded grade. For example, grammar, replicability all have a statistically significant relationship with the target value, which was not the case for the CORD-19 dataset.
3.4 LLM-generated features in classification
@@
In the classification task on respective datasets - binary classification on CORD-19 and ordinal classification on M17+, we tested various models on multiple datasets and chose the best models based on test accuracy. For the purposes of training the machine-learning models, we did not downsample the feature space to only the statistically significant features, as listed in Tables 6, 7 and used all LLM-generated features.
3.4.1 Predictive performance
@@
Table 8 reports three crucial model performance metrics - test accuracy, F1 score, and mean average error (MAE). As the M17+ dataset contains an ordinal target variable, it is also useful to track the MAE, which provides information about the average error of the predicted evaluation. We also report the test F1 scores, which do not differ greatly from the accuracy score, as both datasets were balanced in previous steps.
Reported scores were selected based on the test accuracy of the best models chosen from the models presented in Section 2.5.
@@
For CORD-19, the LLM-feature only set showed significant improvement over the naive baseline model, indicating a 9 percentage point (p.p.) improvement. Furthermore, it scored only 3 p.p. lower than the state-of-the-art SciBERT embeddings for scientific text. The combination of LLM-generated features with TF-IDF terms further improved the test accuracy and F1 score, surpassing SciBERT embeddings and TF-IDF both.
@@
The M17+ features a 5-class ordinal variable, making the task of text classification significantly harder, which is why the overall scores are generally worse for this dataset. Nevertheless, the prediction of human evaluation on the M17+ showed a more pronounced change when compared to the naive baseline model, increasing the test accuracy and F1 score almost twofold. The best-performing model trained on the LLM-features only set was Gradient Boost Classifier and surpassed in terms of test accuracy and F1 score TF-IDF and tied to the state-of-the-art black box AutoGluon model. None of the acquired models reached the results achieved by the model trained on SciBERT embeddings, which scored best in all 3 observed metrics but has the main downside of providing no interpretability to its encoded features.
@@
Table 8: Comparison of Accuracy, F1 Score, and MAE for models trained on each dataset variation, CORD-19 (C19) and M17+.
@@
LLM-features only 0.598 0.370 0.598 0.338 0.402 1.103
Baseline classifier 0.502 0.180 0.502 0.180 0.499 1.620

### v2

@@ 3.1 Computational resources and complexity
This subsection outlines the computational resources and complexity of our feature extraction process. First, we describe local experiments using LLMs on various GPUs—comparing runtime, power consumption, and costs for manually generated features on the CORD-19 and M17+ datasets. Next, we detail our cloud-based approach using the OpenAI Batch API for automatic feature extraction, which leverages batch processing within a 24-hour window for cost efficiency.
3.1.1 LLMs on local GPUs
The non-finetuned Llama2 13B GPTQ Chat model was run locally with the use of GPU computation as the crucial part of the experiment.
For benchmark purposes, we generated the feature ’rigor’ from the CORD-19 dataset on 2 different models utilizing 3 different GPUs. The prompt for ’rigor’ is 126 tokens, while the average abstract contains 229 tokens, averaging 355 input tokens per query. We calculated the costs based on the respective GPU’s worst-case power consumption, the runtime, and the average electricity price of 0.11 USD/kWh. The best performer, time- and money-wise, was Llama2, which was run utilizing the consumer-grade gaming GPU NVIDIA RTX 3090. DeepSeek [^10] https://huggingface.co/deepseek-ai took significantly longer, although it was run on the server GPU NVIDIA L40s, and at higher wattage than Llama2, which was run on entry-level server GPU NVIDIA RTX A4000.
Table 3: Costs of feature extraction using different models and GPUs for determining the rigor feature value based on LLM prompt from Listing 1 for 3,000 instances in CORD-19 dataset
Model@GPU Max Watts Time/1 call Time/3000 calls Price/3000 calls
Llama2@RTX 3090 350W 5.91s 4h 57min \$0.15
Llama2@A4000 140W 11.35s 9h 27min \$0.20
DeepSeek@L40s 300W 63.72s 53h 6min \$1.81
@@ 3.1.1 LLMs on local GPUs
3.1.2 LLMs on cloud GPUs
To facilitate direct comparison and assess the efficiency of cloud-based solutions, the OpenAI Batch API was employed for feature extraction across all datasets (CORD-19, M17+, BANKING77 Hate Speech and Food Hazard). The Batch API allows for the submission of extensive request batches with a guaranteed completion within a user-defined time window; in this case, a 24-hour window was selected. Opting for this 24-hour completion window resulted in a 50% cost reduction compared to synchronous API calls. This cost efficiency is achieved by allowing OpenAI to schedule these tasks during periods of lower demand, optimizing resource utilization. Batches are submitted sequentially.
Table 4 summarizes the costs and batch details associated with the automatic feature extraction process described in Section 2.4.
Table 4: Costs of feature extraction using OpenAI Batch API (model gpt-4o-mini-2024-07-18).
Dataset API Input API Output Total Cost Batches
M17+ \$0.36 \$0.36 \$0.72 3
CORD-19 \$0.47 \$0.48 \$0.96 4
BANKING77 \$1.31 \$1.95 \$3.26 12
Hate Speech \$1.17 \$1.53 \$2.71 11
Food Hazard \$0.66 \$0.76 \$1.42 12
@@ 3.2 LLM-augmented datasets
The features generated by the language model were added to all CORD-19, M17+, BANKING77, Hate Speech, and Food Hazard datasets. Ordinal features such as rigor and novelty for CORD-19 and M17+ were ordinally encoded.
For the user-selected features extracted with local LLM inference, only 2 articles in the CORD-19 dataset were assigned feature values outside of the allowed feature space. These instances were removed, making the final length of CORD-19 2,998 articles (instances). All LLM-generated feature values for the M17+ articles were valid, resulting in a dataset of 2,000 instances. The first few rows from both generated datasets are included in Table 5 and Table 6.
All rows generated in the automatic workflow by gpt-4o-mini with Batch API were valid. A sample of the BANKING77 LLM-generated features can be seen in Table 7.
@@
Table 5: Sample of enriched CORD-19 dataset with added LLM-generated feature values (for user-specified features). Target is a binary variable - with 1 representing works cited more than the median value adjusted in time, 0 represents articles with lower or equal to the median citation count (details in [5]).
@@
Table 6: Sample of enriched M17+ dataset with added LLM-generated feature values (for user-specified features). Target is an ordinal variable of human evaluation of the quality of articles, 1 indicates works of world-class quality, 5 means mediocre articles of low importance and impact.
text contains_question contains_card_mention contains_currency intent_category target
I am still… no yes no activate_my_card 11
What can I… yes yes no activate_my_card 11
Can I link… yes yes no activate_my_card 13
There is a… yes no yes unknown_charge 34
Which fiat… yes yes yes wrong_exchange_rate 36
Table 7: Sample of enriched BANKING77 dataset with added LLM-generated feature values (for automatically generated features). The target variable represents the 77 distinct customer intents.
3.3 Validating LLM features
This subsection documents the validation of LLM-generated features, which was conducted with the help of three techniques. First, the features from each dataset were, along with their dataset’s respective target variable, tested for a statistically significant relationship. This was done on 2,500 bootstrapped samples from the original datasets. Furthermore, Cramér’s V was computed to document the strength of the association of these features and the target of the respective variables. Two selected LLM-generated features were compared against features that were extracted using non-LLM techniques. Finally, we report on a user study aimed at human evaluation of automatically discovered features, which complements the other feature evaluations with a new perspective. While the users may not be able to precisely identify feature quality, they are able to assess whether features subjectively appear as relevant or not.
3.3.1 Analysis of relationship between target and LLM-features
@@
To formally test whether generated features have a relationship with the respective target variables, we employed the chi-squared test of independence at significance level $\alpha=0.05$. We further performed this test on 2,500 bootstrapped samples for each feature to verify the results.
3.3.2 Effect sizes
To validate the results obtained from the chi-squared test and ensure that the statistically significant associations are not due to large sample sizes, we tested the effect size using Cramér’s V [8], see Equation 10.
$V=\sqrt{\frac{\chi^{2}}{n\cdot(k-1)}},$ (10)
where $\chi^{2}$ is the chi-squared statistic, $n$ is the sample size and $k$ is the number of categories in the smaller dimension of the contingency table.
@@ 3.3.2 Effect sizes
Table 8: Sample of statistically significant features according to the relation between target and LLM-generated features (based on user-specified feature list), dataset M17+. The star notation indicates the level of statistical significance: $^{***}p<0.001$, $^{**}p<0.01$, and $^{*}p<0.05$.
Feature Imp. Cramér’s V Feature Imp. Cramér’s V
@@
BioSci [^2] Binary features belonging to disciplines values. *** 0.349 Replicability [^1] Categorical features based on abstract quality and contents. *** 0.322
@@
Basic Medicine [^2] Binary features belonging to disciplines values. *** 0.314 PhysSci [^2] Binary features belonging to disciplines values. *** 0.305
Rigor [^1] Categorical features based on abstract quality and contents. *** 0.296 OtherNatSci [^2] Binary features belonging to disciplines values. *** 0.293
CompSci [^2] Binary features belonging to disciplines values. *** 0.280 OtherMedSci [^2] Binary features belonging to disciplines values. *** 0.265
ChemSci [^2] Binary features belonging to disciplines values. *** 0.261 MedBiotech [^2] Binary features belonging to disciplines values. *** 0.257
AnimalSci [^2] Binary features belonging to disciplines values. *** 0.250 VeterinarySci [^2] Binary features belonging to disciplines values. *** 0.248
Psychology [^2] Binary features belonging to disciplines values. *** 0.246 Novelty [^1] Categorical features based on abstract quality and contents. *** 0.245
EarthSci [^2] Binary features belonging to disciplines values. *** 0.243 OtherAgrolBiotech [^2] Binary features belonging to disciplines values. *** 0.234
@@
Clinical Medicine [^2] Binary features belonging to disciplines values. *** 0.228 Grammar [^1] Categorical features based on abstract quality and contents. *** 0.220
EnviroBiotech [^2] Binary features belonging to disciplines values. *** 0.218 OtherSocSci [^2] Binary features belonging to disciplines values. *** 0.213
HealthSci [^2] Binary features belonging to disciplines values. *** 0.208 Agriculture [^2] Binary features belonging to disciplines values. *** 0.204
Math [^2] Binary features belonging to disciplines values. *** 0.196 IndustBiotech [^2] Binary features belonging to disciplines values. *** 0.191
@@
Mixed methods [^3] Binary features belonging to research values. *** 0.186 LangLiter [^2] Binary features belonging to disciplines values. *** 0.179
We present detailed significance testing results in Table 8 for M17+, which we chose to represent scientometric datasets. Higher values of Cramér’s V are observed for features in the M17+ dataset and target variable. The M17+ dataset has a larger number of features that indicate a statistically significant relationship between the LLM-generated features and the awarded grade. For example, grammar, replicability all have a statistically significant relationship with the target value. Results for Food Hazard, which we chose as the representative of the other datasets, are included in Table 9. The most important features are company_name and hazard_type. The latter is the LLM’s attempt at assessing the hazard_category target variable, which works well. Other features important for hazard category classification are recall_reason, contaminant_type or recall_severity.
Table 9: Sample of statistically significant features according to the relation between hazard-category and LLM-generated features (based on LLM-generated feature list), dataset Food Hazard. The star notation indicates the level of statistical significance: $^{***}p<0.001$, $^{**}p<0.01$, and $^{*}p<0.05$.
Feature Imp. Cramér’s V Feature Imp. Cramér’s V
company_name *** 0.865 hazard_type *** 0.817
recall_reason *** 0.582 contaminant_type *** 0.448
recall_severity *** 0.300 allergen_type *** 0.274
recall_source *** 0.227 product_origin *** 0.204
product_type *** 0.164 distribution_area *** 0.158
recall_frequency *** 0.154 product_packaging *** 0.119
consumer_advice *** 0.116 affected_population *** 0.101
recall_action *** 0.097 recall_notification_method *** 0.088
recall_duration *** 0.087
3.3.3 Validating LLM-features against domain-specific feature generation
To help evaluate the quality of LLM-generated features, we applied non-LLM-based extraction methods to obtain ’grammar’ and ’accessibility’ features for the M17+ dataset. These are then compared with the same features whose values were derived through an LLM.
For ’grammar’, the LanguageTool [16] grammar checker was used through the Python wrapper tool python-language-tool [21].
The LLM-generated feature ’accessibility’ as defined in Table 2 was compared with the SMOG Index [17], which is defined in Equation 11 and is commonly used to evaluate the readability of text and provides the approximate education level required to understand the provided text.
$\text{SMOG Index}=1.043\times\sqrt{\frac{\text{Polysyllabic Words}\times 30}{\text{Number of Sentences}}}+3.1291$ (11)
For the ’accessibility’ feature, we obtained $\rho_{accessibility,SMOGIndex}=0.115$, indicating a weak positive correlation between the features. This goes against our hypothesis that article abstracts marked as more accessible will be awarded a lower SMOG Index. This is likely due to the fact that the SMOG Index uses the number of polysyllabic words per sentence as the measure of readability, while the prompt for the extraction of the ’accessibility’ feature instructed the language model to focus on the clarity of language and how easily the abstracts communicate the main points of the research. Furthermore, the following results were calculated: $\rho_{accessibility,evaluation}=-0.06$ and $\rho_{SMOGIndex,evaluation}=-0.199$, meaning that the SMOG Index correlates with our target variable on this dataset more than the LLM-generated feature ’accessibility’.
In the case of the ’grammar’ feature, the correlation coefficient $\rho=0.141$ indicates a weak correlation between the LLM-generated and LanguageTool assessments. However, when we evaluated the correlations between the target variable and the respective features, we found that the LLM-generated grammar assessment correlates with the target variable with $\rho_{grammar,evaluation}=0.218$ more than the LanguageTool-based feature, which showed no significant correlation with $\rho_{grammar\_ltp,evaluation}=-0.004$.
These findings suggest that while LLM-generated features may not perfectly mirror traditional metrics like the SMOG Index or LanguageTool outputs (showing weak direct correlations), they can capture aspects relevant to the target outcome. This is evidenced by the LLM ’grammar’ feature correlating more strongly with the expert ’evaluation’ score than the standard tool’s output, despite the mixed results for the ’accessibility’ feature.
3.3.4 Predictive performance - LLM features based on user set feature list (scientometric datasets)
@@
In the classification task on respective datasets - binary classification on CORD-19 and ordinal classification on M17+, we tested various models on multiple datasets and chose the best models based on test accuracy. For the purposes of training the machine-learning models, we did not downsample the feature space to only the statistically significant features and used all LLM-generated features.
@@
As the M17+ dataset contains an ordinal target variable, it is also useful to track the mean absolute error (MAE), which provides information about the average error of the predicted evaluation. We also report the test F1 scores, which do not differ greatly from the accuracy score, as both datasets were balanced in previous steps. Table 10 reports three crucial model performance metrics - test accuracy, F1 score, and MAE.
@@
For CORD-19, the LLM-feature only set showed significant improvement over the naive baseline model, indicating a 9 percentage point (p.p.) improvement. Furthermore, it scored only 3 p.p. lower than the state-of-the-art SciBERT embeddings for scientific text. The combination of LLM-generated features with TF-IDF terms further improved the test accuracy and F1 score, surpassing both SciBERT embeddings and TF-IDF.
@@
The M17+ features a 5-class ordinal variable, making the task of text classification significantly harder, which is why the overall scores are generally worse for this dataset. Nevertheless, the prediction of human evaluation on the M17+ showed a more pronounced change when compared to the naive baseline model, increasing the test accuracy and F1 score almost twofold. The best-performing model trained on the LLM-features only set was Gradient Boost Classifier, which surpassed the TF-IDF based model in terms of test accuracy and F1 score and reached results comparable to the state-of-the-art black box AutoGluon model. None of the acquired models reached the results achieved by the model trained on SciBERT embeddings, which scored best in all 3 observed metrics but has the main downside of providing no interpretability to its encoded features.
@@
Table 10: Comparison of Accuracy, F1 Score, and MAE for models trained on each dataset variation, CORD-19 (C19) and M17+. User-selected features only.
@@
LLM-features 0.597 0.355 0.597 0.326 0.403 1.113
Naive classifier 0.502 0.180 0.502 0.180 0.499 1.620
3.3.5 Predictive performance - features selected with LLMs
LLM-generated features that were first selected by an LLM were employed in a classification task and evaluated in the same way as user-selected features in the previous Section 3.3.4.
The choice of the evaluation measures in Tables 11-12 reporting the results depends on the type of the target variable in the dataset. Datasets CORD-19 and M17+ have ordinal/binary targets, while BANKING77, Hate Speech, and Food Hazard datasets have multinomial targets.
The BANKING77 consists of 77 equally represented target classes. The model employing only the LLM-generated features showed an increase in all metrics over the naive uniform classifier, which performed poorly as expected. The same model did not reach the test classification metrics of the TF-IDF-based model. However, combining the LLM-generated features with TF-IDF terms resulted in, albeit a minor, increase in observed metrics over the TF-IDF-only approach.
The model trained on the LLM-feature-only features subset for Hate Speech dataset showed improvement in all metrics besides accuracy when compared to the naïve classifier, which is expected due to class imbalance. Furthermore, while behind in accuracy and precision when compared to the TF-IDF term-based classifier, it showed a 13 p.p. improvement in recall. Upon further inspection, we learned that the class recall of the ’1’ class (hate speech present) increased from 0.14 in the TF-IDF to 0.54 for the LLM-feature only set and to 0.35 for the TF-IDF + LLM-feature set. This result is particularly important as more hate speech comments were successfully detected with LLM-generated features as opposed to the TF-IDF model.
For the Food Hazard dataset, we focused on predicting the ’hazard-category’ target variable, which is one of the tasks of sub-task 1 of the Semeval challenge. Results showed that the LLM-feature only set increased recall over both TF-IDF only and TF-IDF + LLM feature sets while having lower precision. This demonstrates the LLM-feature-based model’s ability to recognize more niche food hazard instances than the traditional TF-IDF-based approach.
Furthermore, the features selected by the LLMs only on the CORD-19, and the M17+ datasets were used in a similar classification experiment to the user-selected features. When directly comparing the selected metric values from Table 10 and Table 11, we can see that both user features and LLM-selected features generated for CORD-19 performed similarly, while user-selected features outperformed the LLM-selected features slightly on the M17+ dataset.
Table 11: Comparison of Accuracy, F1 Score, and MAE for models trained on each dataset variation, CORD-19 (C19) and M17+. Features selected with LLMs only.
Accuracy F1 Score MAE
Model C19 M17+ C19 M17+ C19 M17+
TF-IDF 0.63 0.34 0.62 0.33 0.38 1.07
LLM-features 0.59 0.33 0.59 0.33 0.41 1.17
Naïve classifier 0.50 0.18 0.50 0.18 0.50 1.62
Table 12: Comparison of Accuracy, Recall, and Precision for models trained on each dataset variation: BANKING77 (B77), Hate Speech (Hate), and Food Hazard (Haz). Naïve classifiers were chosen according to the target class characteristics: uniform classifier for BANKING77 and most frequent class for Hate Speech and Food Hazard.
Accuracy Precision Recall
Model B77 Hate Haz B77 Hate Haz B77 Hate Haz
TF-IDF 0.77 0.87 0.91 0.78 0.60 0.80 0.77 0.39 0.55
LLM-features 0.59 0.65 0.93 0.64 0.35 0.77 0.64 0.52 0.64
Naïve classifier 0.01 0.87 0.37 0.01 0.25 0.10 0.01 0.22 0.05